# ViSurf: Visual Supervised-and-Reinforcement Fine-Tuning for Large Vision-and-Language Models

### v1

@@ Abstract
Typical post-training paradigms for Large Vision-and-Language Models (LVLMs) include Supervised Fine-Tuning (SFT) and Reinforcement Learning with Verifiable Rewards (RLVR). SFT leverages external guidance to inject new knowledge, whereas RLVR utilizes internal reinforcement to enhance reasoning capabilities and overall performance. However, our analysis reveals that SFT often leads to sub-optimal performance, while RLVR struggles with tasks that exceed the model’s internal knowledge base. To address these limitations, we propose ViSurf (Visual Supervised-and-Reinforcement Fine-Tuning), a unified post-training paradigm that integrates the strengths of both SFT and RLVR within a single stage. We analyze the derivation of the SFT and RLVR objectives to establish the ViSurf objective, providing a unified perspective on these two paradigms. The core of ViSurf involves injecting ground-truth labels into the RLVR rollouts, thereby providing simultaneous external supervision and internal reinforcement. Furthermore, we introduce three novel reward control strategies to stabilize and optimize the training process. Extensive experiments across several diverse benchmarks demonstrate the effectiveness of ViSurf, outperforming both individual SFT, RLVR, and two-stage SFT →RLVR. In-depth analysis corroborates these findings, validating the derivation and design principles of ViSurf.
@@ 1 Introduction
Developing Large Vision-and-Language Models (LVLMs) that excel in diverse visual perception tasks is a promising direction for visual intelligence. To this end, prior work has predominantly relied on two training paradigms: Supervised Fine-Tuning (SFT) [17, 38, 1] and Reinforcement Learning with Verifiable Rewards (RLVR) [20, 21].
@@
Both these two paradigms have its advantages and disadvantages. SFT directly optimizes the model using expert-annotated data. This approach provides explicit external guidance, enabling the model to memorize the target distribution. However, it often exhibits limited generalization capabilities and can lead to catastrophic forgetting of pre-trained knowledge. Recently, RLVR has garnered significant research attention. Methods such as Group Relative Policy Optimization (GRPO) [34] or Dynamic Sampling Policy Optimization (DAPO) [42] are on-policy algorithms wherein an initial policy generates rollouts that are evaluated by pre-defined reward functions, and the policy is subsequently optimized based on this internal feedback. By leveraging internal reinforcement signals, RLVR mitigates catastrophic forgetting and often achieves superior generalization. Nevertheless, its performance can degrade when tasks extend beyond the initial model’s knowledge base. This limitation is particularly pronounced in LVLMs compared to Large Language Models (LLMs), as the visual and textual domains exhibit significant variance in both input modalities and expected outputs. We evaluate these two paradigms across a diverse set of vision-and-language tasks, with our key findings summarized in Figure 1. The results indicate that SFT is more effective for tasks that fall outside the pre-training distribution of LVLMs, whereas RLVR yields superior performance on tasks that align with its pre-existing knowledge base. We also provide a case study in Section 3.2 to illustrate this phnomenon. Although a sequential SFT →RLVR pipeline leverages their complementary strengths, it incurs the combined computational cost of both stages and remains susceptible to catastrophic forgetting during the initial SFT phase.
To address the aforementioned limitations, we propose ViSurf (Visual Supervised-and-Reinforcement Fine-Tuning), a unified, single-stage post-training paradigm designed to integrate the complementary advantages of SFT and RLVR. We begin with an analysis of the underlying objectives and gradients of SFT and RLVR. We theoretically demonstrate that their formulations share similar patterns, enabling them to be integrated into a single, unified objective—the ViSurf objective. While subtle differences exist, the gradient of ViSurf objective can be interpreted as a composite of the gradients from both SFT and RLVR. Based on the theoretical analysis, we design three reward control strategies to stabilize training. Specifically, for the ground-truth labels, we i) aligning them with rollouts preference, ii) eliminating thinking reward for them, and iii) smoothing the reward for them. The implementation of ViSurf is consequently simplified to interleaving ground-truth demonstrations with on-policy rollouts in a unified training phase.
@@
We conduct extensive experiments across various domains, with a comparative summary presented in Figure 2. The results indicate that our method, ViSurf, achieves superior performance than SFT, RLVR and SFT →RLVR. Models trained with ViSurf also exhibit reasoning abilities similar to those established in previous works [20, 21]. Furthermore, ViSurf successfully mitigates catastrophic forgetting, as evidenced by its stable performance on VQA tasks. The ablation study confirms the critical contribution of the proposed reward control mechanism. Additional in-depth analysis provides empirical validation for our theoretical analysis and offers insights into the operational principles of ViSurf. Our contributions are summarized as follows:
@@
  • Empirically, ViSurf outperforms existing methods (SFT, RLVR, and SFT →RLVR). Additional analyses offer a thorough understanding of its operational mechanics.
@@ 2.1 Supervised Fine-tuning for LVLMs
Supervised Fine-Tuning (SFT) is a predominant paradigm for training Large Vision-Language Models (LVLMs). This approach involves fine-tuning pre-trained models on expert-annotated data. The pioneering work in this area is LLaVA [17], which has inspired numerous subsequent models. Notable examples include the LLaVA-series [13, 18], QwenVL-series [38, 1], MGM-series [14, 43, 36], InternVL [3] and LLaMA-3.2 [27], all of which adopt this paradigm. SFT has proven particularly effective for adapting LVLMs to diverse downstream applications, such as image quality assessment [41], visual counting [5], and autonomous driving [40].
@@ 2.2 Reinforcement Learning for LVLMs
Reinforcement Learning (RL) is a standard method for fine-tuning Large Vision-Language Models (LVLMs). Among RL algorithms, Direct Preference Optimization (DPO) [30] relies on pre-collected human preference datasets, which can be costly to produce. Similarly, Proximal Policy Optimization (PPO) [32] requires a well-trained reward model to evaluate responses generated by the policy. Recently, RLVR algorithms, such as GRPO [34] and DAPO [42], have gained attention for their ability to assess model outputs against objective criteria. This approach reduces the dependency on manually annotated data and pre-trained reward models. The effectiveness of RLVR for LVLMs has been demonstrated in recent works [20, 22, 21, 19, 9] such as SegZero [20] and VisualRFT [22].
@@ 3.1 Preliminary
Supervised Fine-Tuning (SFT) optimizes $\pi_{\theta}$ against a set of pre-annotated ground-truth labels, $\mathcal{D}_{\text{label}}=\{y_{i}\}^{N}_{i=1}$. The objective is to minimize the negative log-likelihood of the labels:
@@
where $y$ corresponds to $(v,t)$. A more precise notation would be $(v,t,{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y})\sim\text{zip}(\mathcal{D}_{\text{input}},{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\mathcal{D}_{\text{label}}})$. Nevertheless, we retain the current notation, $(v,t)\sim\mathcal{D}_{\text{input}},{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}$, for the sake of clarity and ease of comparison in the subsequent discussion.
Reinforcement Learning with Verifiable Rewards (RLVR). We illustrate RLVR using the on-policy Group Relative Policy Optimization (GRPO) algorithm [34]. GRPO optimizes the policy $\pi_{\theta}$ using a verifiable reward function, which typically combines measures of output format and accuracy [7, 21, 20]. For a given input $(v_{i},t_{i})\in\mathcal{D}_{\text{input}}$, the old policy $\pi_{\theta_{old}}$ (from a previous optimization step) generates a group of $G$ rollouts $\{o_{j}\}^{G}_{j=1}$ by sampling with different random seeds. Each rollout $o_{j}$ is then evaluated by a reward function $r(\cdot)$, resulting in a set of rewards $\{r(o_{j})\}^{G}_{j=1}$. The advantage for each rollout is subsequently computed as follows:
@@
The objective of RLVR is to minimize the following equation:
@@
$\hskip 17.00024pt\hskip 17.00024pt\left.\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)},\;1-\epsilon,\;1+\epsilon\Biggr)\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\Bigg\}\right].$
For simplicity, both in equation and in our practical implementation, we omit the KL divergence term.
@@ 3.2 Case Study: Non-Object Scenarios
We present a case study on non-object referring expression segmentation. In this task, instructions comprise both correct expressions (referring to existing objects) and incorrect ones (where the referred objects are absent). For this study, we utilize the VisionReasoner model [21], which is initialized with Qwen2.5VL [1] and SAM2 [31] and has demonstrated strong performance on standard referring segmentation tasks. Figure 3(a) illustrates the non-object segmentation and the architecture of VisionReasoner. Detailed experimental settings are demonstrated Section 4.1.
@@
The gRefCOCO [16] serves as the benchmark for this scenario. We compare gIoU (average IoU across images) and N-Acc (accuracy in identifying non-existent object). Quantitative and qualitative results are presented in Figure 3(b)-(c). Our analysis reveals that SFT yields sub-optimal gIoU and fails to produce a coherent reasoning process, while models trained with RLVR tend to generate a mask even when no relevant object is detected. The latter issue arises because the model, relying solely on self-rollouts, cannot correct itself to produce a ‘no object’ output. In essence, RLVR enhances overall performance through internal self-guidance, whereas SFT provides crucial external guidance when self-rollouts fail. This analysis leads to a key question: how can we integrate the benefits of both training paradigms in a single training stage efficiently?
@@ 3.3 Combining SFT and RLVR
To harness the complementary benefits of SFT and RLVR, we propose ViSurf: a unified, single-stage post-training algorithm. This section elaborates on the derivation of the ViSurf algorithm.
@@
The objective of ViSurf is to minimize the following equation:
@@
Gradient Analysis of ViSurf. The gradient of Equation 8 can be derived using approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ and log-derivative trick. We also omit clip operation for simplicity:
@@
To better illustrate the structure of the gradient, we reformulate Equation 9 as following:
@@
Relation to SFT and RLVR. We analyze the gradients of ViSurf, SFT, and RLVR to explain their connections. The RLVR term in Equation 10 is structurally identical to the standard RLVR gradient in Equation 5, differing only in its scaling coefficient ($\frac{1}{G+1}\hat{A}_{j}$ vs. $\frac{1}{G}\hat{A}_{j}$). Similarly, the SFT term in Equation 10 resembles the SFT gradient from Equation 4, with two key distinctions: (i) the coefficient is weighted by $\frac{1}{G+1}\hat{A}_{y}$ instead of 1, and (ii) the use of the approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$. This approximation implies that the ground-truth label $y$ should align with the model’s internal generative preference to be effective. Crucially, Equation 10 integrates both the external guidance from SFT and the internal guidance from RLVR into a unified gradient.
@@ 3.4 Reward Control for Ground-Truth Label
Eliminating Thinking Reward for Ground-truth Labels. Since the ground-truth labels lack annotated reasoning path, we assign a reasoning format score of zero to them. This operation ensures the model learns reasoning trace directly from its self-rollouts without being biased by missing external annotations.
Smoothing the Reward for Ground-truth Labels. Prior to advantage estimation, we compare the maximum reward among generated rollouts, $\max\{\{\text{r}(o_{j})\}_{j=1}^{G}\}$, against the ground-truth reward $\text{r}(y)$. If $\max\{\{\text{r}(o_{j})\}_{j=1}^{G}\}\geq\text{r}(y)$, it indicates the policy model $\pi_{\theta}$ has already produced a high-quality output without external guidance. In this case, we set $\text{r}(y)=\text{mean}\{\{\text{r}(o_{j})\}_{j=1}^{G}\}$. This smoothing ensures that the advantage for the ground-truth, $\hat{A}_{y}$, becomes zero (as per Equation 7), effectively eliminating the external supervision signal when it is unnecessary.

### v2

@@ Abstract
Typical post-training paradigms for Large Vision-and-Language Models (LVLMs) include Supervised Fine-Tuning (SFT) and Reinforcement Learning with Verifiable Rewards (RLVR). SFT leverages external guidance to inject new knowledge, whereas RLVR utilizes internal reinforcement to enhance reasoning capabilities and overall performance. However, our analysis reveals that SFT often leads to sub-optimal performance, while RLVR struggles with tasks that exceed the model’s internal knowledge base. To address these limitations, we propose ViSurf (Visual Supervised-and-Reinforcement Fine-Tuning), a unified post-training paradigm that integrates the strengths of both SFT and RLVR within a single stage. We analyze the derivation of the SFT and RLVR objectives to establish the ViSurf objective, providing a unified perspective on these two paradigms. The core of ViSurf involves injecting ground-truth labels into the RLVR rollouts, thereby providing simultaneous external supervision and internal reinforcement. Furthermore, we introduce three novel reward control strategies to stabilize and optimize the training process. Extensive experiments across several diverse benchmarks demonstrate the effectiveness of ViSurf, outperforming both SFT, RLVR, and two-stage SFT → RLVR. In-depth analysis corroborates these findings, validating the derivation and design principles of ViSurf. The code will be publicly available.
@@ 1 Introduction
Developing Large Vision-and-Language Models (LVLMs) that excel in diverse visual perception tasks is a promising direction for visual intelligence. To this end, prior work has predominantly relied on two training paradigms: Supervised Fine-Tuning (SFT) [liu2023visual, wang2024qwen2vl, bai2025qwen25vl] and Reinforcement Learning with Verifiable Rewards (RLVR) [liu2025seg, liu2025visionreasoner].
@@
Every coin has two sides, and these paradigms are no exception. SFT directly optimizes the model using expert-annotated data. This approach provides explicit external guidance, enabling the model to memorize the target distribution. However, it often exhibits sub-optimal performance and can lead to catastrophic forgetting of pre-trained knowledge. Recently, RLVR has garnered significant research attention. The on-policy RLVR methods start with an initial policy and iteratively generates a group of rollouts that are evaluated by pre-defined reward functions, then the policy is subsequently optimized based on this internal feedback, such as Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] and Dynamic Sampling Policy Optimization (DAPO) [yu2025dapo]. By leveraging internal reinforcement signals, RLVR mitigates catastrophic forgetting and often achieves superior performance. Nevertheless, its performance can degrade when tasks extend beyond the initial model’s knowledge base. We evaluate these two paradigms across a diverse set of vision-and-language tasks, with our key findings summarized in Figure 1. The results indicate that SFT is more effective for tasks that fall outside the pre-training distribution of LVLMs, whereas RLVR yields superior performance on tasks that align with its pre-existing knowledge base. We also provide a case study in Section 3.2 to illustrate this phnomenon. Although a sequential SFT → RLVR pipeline tries to leverage their complementary strengths, this two-stage method incurs the combined computational cost of both stages and remains susceptible to catastrophic forgetting during the initial SFT phase.
To address the aforementioned limitations, we propose ViSurf (Visual Supervised-and-Reinforcement Fine-Tuning), a unified, single-stage post-training paradigm designed to integrate the complementary advantages of SFT and RLVR. We begin with an analysis of the underlying objectives and gradients of SFT and RLVR. We theoretically demonstrate that their gradients share similar patterns, enabling them to be integrated into a single, unified objective, the ViSurf objective. Unlike prior methods [zhang2025policy, ma2025learning] that combine SFT and RLVR through simple objective addition, ViSurf offers a unified perspective for these two training paradigms. While subtle differences exist, the gradient of ViSurf objective can be interpreted as a composite of the gradients from both SFT and RLVR. Based on the theoretical analysis, we design three reward control strategies to stabilize training. Specifically, for the ground-truth labels, we i) aligning them with rollouts preference, ii) eliminating thinking reward for them, and iii) smoothing the reward for them. The implementation of ViSurf is consequently simplified to interleaving ground-truth demonstrations with on-policy rollouts in a unified training phase.
@@
We conduct extensive experiments across various domains, with a comparative summary presented in Figure 2. The results indicate that our method, ViSurf, achieves superior performance than SFT, RLVR and SFT → RLVR. Models trained with ViSurf also exhibit reasoning abilities similar to those established in previous works [liu2025seg, liu2025visionreasoner]. Furthermore, ViSurf successfully mitigates catastrophic forgetting, as evidenced by its stable performance on VQA tasks. The ablation study confirms the critical contribution of the proposed reward control mechanism. Additional in-depth analysis provides empirical validation for our theoretical analysis and offers insights into the operational principles of ViSurf. Our contributions are summarized as follows:
@@
  • Empirically, ViSurf outperforms existing post-training methods (SFT, RLVR, and SFT → RLVR). Additional analyses offer a thorough understanding of its mechanics.
@@ 2.1 Supervised Fine-tuning for LVLMs
Supervised Fine-Tuning (SFT) is a predominant paradigm for training Large Vision-Language Models (LVLMs). This approach involves fine-tuning pre-trained models on expert-annotated data. The pioneering work in this area is LLaVA [liu2023visual], which has inspired numerous subsequent models. Notable examples include the LLaVA-series [li2024llavaov, liu2024improved], QwenVL-series [wang2024qwen2vl, bai2025qwen25vl], MGM-series [li2024mini, zhong2024lyra, wang2025mgm], and InternVL [chen2024internvl], all of which adopt this paradigm. SFT has proven particularly effective for adapting LVLMs to diverse downstream applications, such as image quality assessment [you2025teaching], visual counting [deitke2024molmo], and autonomous driving [xu2025drivegpt4].
@@ 2.2 Reinforcement Learning for LVLMs
Reinforcement Learning (RL) is a standard method for fine-tuning Large Vision-Language Models (LVLMs). Among RL algorithms, Direct Preference Optimization (DPO) [rafailov2023direct] relies on pre-collected human preference datasets, which can be costly to produce. Similarly, Proximal Policy Optimization (PPO) [schulman2017proximal] requires a well-trained reward model to evaluate responses generated by the policy. Recently, RLVR algorithms, such as GRPO [shao2024deepseekmath] and DAPO [yu2025dapo], have gained attention for their ability to assess model outputs against objective criteria. This approach reduces the dependency on manually annotated data and pre-trained reward models. The effectiveness of RLVR for LVLMs has been demonstrated in recent works [liu2025seg, liu2025visual, liu2025visionreasoner, liu2025empowering, huang2025vision] such as SegZero [liu2025seg] and VisualRFT [liu2025visual].
@@ 3.1 Preliminary
Supervised Fine-Tuning (SFT) optimizes $\pi_{\theta}$ against a set of ground-truth labels, $\mathcal{D}_{\text{label}}=\{y_{i}\}^{N}_{i=1}$. The objective is to minimize the negative log-likelihood of the labels:
@@
where $y$ corresponds to $(v,t)$. A more precise notation would be $(v,t,{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y})\sim\text{zip}(\mathcal{D}_{\text{input}},{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\mathcal{D}_{\text{label}}})$. Nevertheless, we retain the current notation, $(v,t)\sim\mathcal{D}_{\text{input}},{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}$, for clarity and ease of comparison in the subsequent discussion.
Reinforcement Learning with Verifiable Rewards (RLVR). We illustrate RLVR using the on-policy Group Relative Policy Optimization (GRPO) algorithm [shao2024deepseekmath]. GRPO optimizes the policy $\pi_{\theta}$ using a verifiable reward function, which typically combines measures of output format and accuracy [guo2025deepseek, liu2025visionreasoner, liu2025seg]. For a given input $(v_{i},t_{i})\in\mathcal{D}_{\text{input}}$, the old policy $\pi_{\theta_{old}}$ (from a previous optimization step) generates a group of $G$ rollouts $\{o_{j}\}^{G}_{j=1}$ by sampling with different random seeds. Each rollout $o_{j}$ is then evaluated by a reward function $r(\cdot)$, resulting in a set of rewards $\{r(o_{j})\}^{G}_{j=1}$. The advantage for each rollout is subsequently computed as follows:
@@
The objective of RLVR is to minimize the equation:
@@
$\hskip 17.00024pt\hskip 17.00024pt\left.\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)},\;1-\epsilon,\;1+\epsilon\Biggr)\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\Bigg\}\right],$
where the $\epsilon$ is a constant that controls the clipping boundary. For simplicity, both in equation and in our practical implementation, we omit the KL divergence term.
@@ 3.2 Case Study: Non-Object Scenarios
We present a case study on non-object referring expression segmentation. In this task, instructions comprise both correct expressions (referring to existing objects) and incorrect ones (where the referred objects are absent). For this study, we utilize the VisionReasoner model [liu2025visionreasoner], which is initialized with Qwen2.5VL [bai2025qwen25vl] and SAM2 [ravi2024sam2] and has demonstrated strong performance on standard referring segmentation tasks. Figure 3(a) illustrates the non-object segmentation and the architecture of VisionReasoner. Detailed experimental settings are demonstrated Section 4.1.
@@
The gRefCOCO [liu2023gres] serves as the benchmark for this scenario. We compare gIoU (average IoU across images) and N-Acc (accuracy in identifying non-existent object). Quantitative and qualitative results are presented in Figure 3(b)-(c). Our analysis reveals that model trained by SFT achieves suboptimal gIoU performance but tends to learn to correctly identify the absence of objects. In contrast, model trained with RLVR exhibits a more coherent reasoning process and attain higher overall gIoU scores, yet they often generate object masks even when no relevant object is present. This limitation arises because the RLVR model, relying solely on self-rollouts, lacks the corrective mechanism necessary to produce a “no object” output. In essence, RLVR enhances overall performance through internal self-guidance, whereas SFT provides crucial external guidance when self-rollouts fail. This analysis leads to a key question: how can we integrate the benefits of both training paradigms in a single training stage efficiently?
@@ 3.3 Combining SFT and RLVR
To harness the complementary benefits of SFT and RLVR, we propose ViSurf: a unified, single-stage post-training algorithm. This section elaborates on the derivation of the ViSurf algorithm and its relation to SFT and RLVR.
@@
The objective of ViSurf is to minimize the equation:
@@
Gradient Analysis of ViSurf. The gradient of Equation 8 can be derived using approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ and log-derivative trick. We omit clip operation for simplicity:
@@
Relation to SFT and RLVR. We analyze the gradients of ViSurf, SFT, and RLVR to explain their connections. To better illustrate the structure of the gradient, we reformulate Equation 9 as following:
@@
The RLVR term in Equation 10 is structurally identical to the standard RLVR gradient in Equation 5, differing only in its scaling coefficient ($\frac{1}{G+1}\hat{A}_{j}$ vs. $\frac{1}{G}\hat{A}_{j}$). Similarly, the SFT term in Equation 10 resembles the SFT gradient from Equation 4, with two key distinctions: (i) the coefficient is weighted by $\frac{1}{G+1}\hat{A}_{y}$ instead of 1, and (ii) the use of the approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$. This approximation implies that the ground-truth label $y$ should align with the model’s internal generative preference to be effective. Crucially, Equation 9 integrates both the external guidance from SFT and the internal guidance from RLVR.
@@ 3.4 Reward Control for Ground-Truth Label
Eliminating Thinking Reward for Ground-truth Labels. Since the ground-truth labels lack annotated reasoning path, we assign a reasoning format score of zero to them. This operation ensures the model learns reasoning trace directly from its self-rollouts without being biased by missing external reasoning annotations.
Smoothing the Reward for Ground-truth Labels. Prior to advantage estimation, we compare the maximum reward among generated rollouts, $\max\{\{\text{r}(o_{j})\}_{j=1}^{G}\}$, against the ground-truth reward $\text{r}(y)$. If $\max\{\{\text{r}(o_{j})\}_{j=1}^{G}\}\geq\text{r}(y)$, it indicates the policy model $\pi_{\theta}$ has already produced a high-quality output without external guidance. In this case, we set $\text{r}(y)=\text{mean}\{\{\text{r}(o_{j})\}_{j=1}^{G}\}$. This smoothing ensures that the advantage for the ground-truth, $\hat{A}_{y}$, becomes zero (as per Equation 7), eliminating the external supervision signal when it is unnecessary.